# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 10 Artin groups with complete bipartite Dynkin diagrams
Suppose $\Lambda=\Lambda_{m,n}$ which is the join of $m$ vertices and $n$ vertices. The case that one of $m$ and $n$ is 1 follows from Lemma 10.5. Now we assume $m>1$ and $n>1$. Let $f:\Lambda\to C$ be as in Lemma 10.6. We claim whenever $f$ is a composition of two special foldings $f^{\prime}:\Lambda\to\Lambda^{\prime}$ and $f^{\prime\prime}:\Lambda^{\prime}\to C$ with $\Lambda^{\prime}$ being complete bipartite, then $\Delta_{\Lambda,f^{\prime}}$ is contractible. The lemma follows from this claim by taking $f^{\prime\prime}$ to be identity.
@@
such that $f=f_{1}\circ f_{2}\circ\cdots\circ f_{n}$, each $\Lambda_{i}$ is complete bipartite and $\Lambda_{i+1}$ has exactly one vertex less than $\Lambda_{i}$. Let $g_{i}=f_{1}\circ\cdots\circ f_{i}$ and $\Delta_{i}=\Delta_{\Lambda,g_{i}}$. The sequence can be arranged such that $\Delta_{\Lambda,f^{\prime}}=\Delta_{i}$ for some $i$. So it suffices to show $\Delta_{i}$ is contractible for any $i$. We set $\Delta_{0}=\Delta_{\Lambda}$.
For each $i$, let $v_{i}\in\Lambda_{i}$ be the unique vertex such that $f^{-1}_{i}(v_{i})$ has more than one vertices. For each edge in $\Delta_{i}$ whose vertices are of type $f^{-1}_{i+1}(v_{i+1})$, we add a new vertex in this edge which is the midpoint of this edge, and say this new vertex has type $\hat{m}_{i}$. Cut each top dimensional simplex in $\Delta_{i}$ along the new vertex into two simplices, and let the resulting complex by $\Delta^{\prime}_{i}$. Then there is a natural embedding $\iota_{i+1}:\Delta_{i+1}\to\Delta_{i}$ mapping vertices of type $\hat{v}$ to vertices of type $\hat{f}^{-1}_{i+1}(v)$ for $v\in\Lambda_{i+1}\setminus\{v_{i+1}\}$, and mapping vertices of type $\hat{v}_{i+1}$ to vertices of type $\hat{m}_{i}$. The image of $\iota_{i+1}$ is the full subcomplex of $\Delta_{i}$ spanned by vertices whose types are either $\hat{m}_{i}$ or inside $f^{-1}_{i+1}(\Lambda_{i+1}\setminus\{v_{i+1}\})$.
By Lemma 10.6, Theorem 3.28 and [Hua23, Lemma 10.3], we know $\Delta_{n}$ is contractible. Next we will show $\Delta_{i}$ and $\Delta_{i+1}$ are homotopic equivalent for $10\leq i\leq n-$ 1. By the description of how $\Delta_{i+1}$ sits as a subcomplex of $\Delta^{\prime}_{i}$ via $\iota_{i+1}$ in the previous paragraph, it suffices to show for any $v\in f^{-1}_{i+1}(v_{i+1})$ and any vertex $x\in\Delta_{i}$ of type $\hat{v}$, $\operatorname{lk}(x,\Delta_{i})$ is contractible, as this would imply $\Delta^{\prime}_{i}$ deformation retracts onto $\Delta_{i+1}$. As $\Lambda_{i}$ is complete bipartite, $f^{-1}_{i+1}(v_{i+1})$ is contained in a join factor of $\Lambda^{\prime}_{i}$ of $\Lambda_{i}$. As $f^{-1}_{i+1}(v_{i+1})$ has two elements, we know $\Lambda_{i}\setminus\{v\}$ is connected. Let $\Lambda_{i,v}$ be the unique component of $\Lambda_{i}\setminus\{v\}$. Then $\Lambda_{i,v}$ is also complete bipartite. Let $\Theta=g^{-1}_{i}(\Lambda_{i,v})$. By [Hua23, Lemma 10.4],
@@
Let $\Lambda$ be a complete bipartite Dynkin diagram. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
@@
Note that when $\Lambda$ has $\leq 4$ vertices, $A_{\Lambda}$ satisfies $K(\pi,1)$-conjecture by Theorem 9.2. Now the theorem follows by induction on the number of vertices in $\Lambda$, using Proposition 10.2, Lemma 10.7 and Theorem 3.1. ∎
@@
Let $\Lambda$ be a tree Dynkin diagram with a collection of open edges $E$ with label $\geq 6$ such that each component of $\Lambda\setminus(\cup_{e\in E}\{e\})$ is either spherical or a star. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
@@
By Theorem 1.1 and [Hua23, Proposition 9.12], it suffices to show $\Delta_{\Lambda}$ satisfies labeled 4-wheel condition whenever $\Lambda$ is a star. Take a maximal linear subgraph $\Lambda^{\prime}\subset\Lambda$. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is bowtie free - the case when $\Lambda$ has $\leq 3$ vertices follows from Theorem 3.17 and Lemma 10.3, the case when $\Lambda$ has 4 vertices follows from Proposition 10.2 and the case when $\Lambda$ has $\geq 5$ vertices follows from Lemma 10.5. Then Lemma 3.20 implies that $\Delta_{\Lambda}$ satisfies labeled 4-wheel condition. ∎
@@ 11 $K(\pi,1)$ for some higher-dimensional families
Let $\Lambda$ be a connected Dynkin diagram with an induced sub-diagram $\Lambda^{\prime}\subset\Lambda$ such that
@@
    $\Lambda^{\prime}$ is the Dynkin diagram of a 3-dimensional irreducible affine Coxeter group;
@@
    for any vertex $s\in\Lambda^{\prime}$, each component of $\Lambda\setminus\{s\}$ is either spherical, or has type in $\{\widetilde{A}_{3},\widetilde{B}_{3},\widetilde{C}_{n}\}$.
@@
Let $\mathcal{C}$ be the class of Dynkin diagrams satisfying the assumptions of Proposition 11.1. Let $\Lambda_{0}$ be a 5-cycle with its consecutive edges labeled by $\{3,3,4,3,4\}$. Let $\mathcal{C}_{T}$ be the subclass made of members of $\mathcal{C}$ that are trees. Let $\mathcal{C}^{\prime}$ be the collection of all $\Lambda\in\mathcal{C}$ such that there exists $\Lambda^{\prime}\subset\Lambda$ of type $\widetilde{A}_{3}$ with Assumption 2 satisfied. We first show that element of $\mathcal{C}=\mathcal{C}_{T}\sqcup\mathcal{C}^{\prime}\sqcup\{\Lambda_{0}\}$. Indeed, take $\Lambda\in\mathcal{C}$ and let $\Lambda^{\prime}\subset\Lambda$ be as in Assumption 2. First we consider the case $\Lambda^{\prime}$ has type $\widetilde{C}_{3}$. Let $\{s_{i}\}_{i=1}^{4}$ be consecutive vertices in $\Lambda^{\prime}$. By our assumption and the classification of spherical and Euclidean Dynkin diagrams, for $i=1,4$, the only possibility types for each component of $\Lambda\setminus\{s_{i}\}$ are $\{B_{n},F_{4},\widetilde{B}_{3},\widetilde{C}_{n}\}$. If $\Lambda$ is not a tree, then it must contain an embedded cycle $C$. As each component of $\Lambda\setminus\{s_{i}\}$ is a tree for $i=1,4$, we know $\{s_{1},s_{4}\}\subset C$. Note that $s_{2}\in C$, otherwise $C\cup\overline{s_{3}s_{4}}$ is contained in a component of $\Lambda\setminus\{s_{2}\}$, which is not possible by our assumption (as the label of $\overline{s_{3}s_{4}}$ is $4$). Similarly $s_{3}\in C$. Thus $\Lambda^{\prime}\subset C$. Let $e_{i}$ be the edge of $C$ that is outside $\Lambda^{\prime}$ and contains $s_{i}$ for $i=1,4$. As $\Lambda^{\prime}$ is an induced subgraph of $\Lambda$, $s_{1}\notin e_{4}$ and $s_{4}\notin e_{1}$. Thus $e_{1}\cup\overline{s_{1}s_{2}}\cup\overline{s_{2}s_{3}}$ is contained in a component of $\Lambda\setminus\{s_{4}\}$, and by our assumption, the only possibility of this component is $F_{4}$. Similarly, the component of $\Lambda\setminus\{s_{1}\}$ containing $\overline{s_{2}s_{3}}\cup\overline{s_{3}s_{4}}\cup e_{4}$ is of type $F_{4}$. It follows from Assumption 2 of the proposition that $\Lambda=\Lambda_{0}$. Now we consider the case that $\Lambda^{\prime}$ has type $\widetilde{B}_{3}$. Let $a$ be the center vertex and $\{b_{i}\}_{i=1}^{3}$ be leave vertices of $\Lambda^{\prime}$ with $m_{a,b_{1}}=4$. For $1\leq i\leq 3$, let $\Lambda_{i}$ be the component of $\Lambda\setminus\{b_{i}\}$ containing $a$. Then for $i=2,3$, $\Lambda_{i}$ contains an edge labeled by 4, hence is a tree by the classification of spherical and Euclidean Dynkin diagrams. If $\Lambda_{1}$ is a tree, then $\Lambda$ is a tree. Otherwise $\Lambda_{1}$ is of type $\widetilde{A}_{3}$. Then for $i=2,3$, $\Lambda_{i}$ contains a path of length 3, with its edges labeled by $4,3,3$. Thus $\Lambda_{i}$ is of type $B_{n}$ or $\widetilde{C}_{n}$ for $i=2,3$. However, only type $B_{n}$ and $n=4$ is possible, otherwise $\Lambda_{1}$ is not of type $\widetilde{A}_{3}$. This implies that $\Lambda$ is obtained from a diagram of type $\widetilde{A}_{3}$ by adding an extra edge of label $4$. Thus $\Lambda\in\mathcal{C}^{\prime}$.
@@
Next we show $\mathcal{C}^{\prime}$ satisfy the assumptions of [Hua23, Corollary 7.3] with $\mathcal{C}_{2}=\mathcal{C}^{\prime}$, and $\mathcal{C}_{1}$ be the class containing only the diagram of type $\widetilde{A}_{3}$, hence each element in $\mathcal{C}^{\prime}$ satisfies the $K(\pi,1)$-conjecture. Note that Assumptions $2,3,4$ of [Hua23, Corollary 7.3] are clear. It remains to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible. For vertex $s\in\Lambda^{\prime}$, Let $\Lambda_{s}$ be the component of $\Lambda\setminus\{s\}$ containing the rest of the vertices of $\Lambda^{\prime}$. If $\Lambda_{s}$ is not a tree for some $s$, then $\Lambda_{s}$ is of type $\widetilde{A}_{3}$. Then $\Lambda$ is a complete bipartite graph $K_{2,3}$. By [Hua23, Theorem 10.7 and Corollary 10.10], $\Delta_{\Lambda}$ is contractible. Hence $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible by Lemma 3.5. Now we assume $\Lambda_{s}$ is a tree for each $s$. By Theorem 3.28 and Lemma 3.3, it suffices to show $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. However, this follows from Theorem 3.17 if $\Lambda_{s}$ is spherical, and Corollary 5.2 if $\Lambda_{s}$ is of type $\widetilde{C}_{n}$. If $\Lambda_{s}$ is of type $\widetilde{B}_{3}$, then by Theorem 3.22 and Theorem 5.6, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 3.35. Thus by Proposition 8.1 $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free.
@@
As the $K(\pi,1)$-conjecture for Artin group with diagram $\Lambda_{0}$ follows from [Hua23, Theorem 10.9], it remains to show diagrams in $\mathcal{C}_{T}$ satisfies $K(\pi,1)$-conjecture. We will show $\mathcal{C}_{T}$ satisfy the assumptions of [Hua23, Corollary 7.3] with $\mathcal{C}_{2}=\mathcal{C}_{T}$ and $\mathcal{C}_{1}=\{\widetilde{C}_{3},\widetilde{B}_{3}\}$. Again Assumptions $2,3,4$ of [Hua23, Corollary 7.3] are clear, and it suffices to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible with $\Lambda^{\prime}\in\mathcal{C}_{1}$ and $\Lambda\in\mathcal{C}_{2}$.
@@
Case 1: $\Lambda^{\prime}$ is of type $\widetilde{C}_{3}$. Let consecutive vertices of $\Lambda^{\prime}$ be $\{s_{i}\}_{i=1}^{4}$. Let $\Lambda_{1}$ be the component of $\Lambda\setminus\{s_{1}\}$ containing $\Lambda^{\prime}\setminus\{s_{1}\}$. By Assumption 2, $\Lambda_{1}$ is either spherical or irreducible 3-dimensional Euclidean. We claim the vertex set of $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$, endowed with the order induced from $s_{2}<s_{3}<s_{4}$ is a bowtie free and upward flag poset. As $\Lambda_{1}$ is a tree, $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ is indeed a poset. If $\Lambda_{1}$ is spherical, then either $\Lambda_{1}$ is of type $B_{n}$ or type $F_{4}$. Then $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ being bowtie free and upward flag follows from Theorem 3.17, Theorem 3.22 and Proposition 4.2. If $\Lambda_{1}$ is of type $\widetilde{C}_{n}$, then the claim follows from Corollary 5.2. If $\Lambda_{1}$ is of type $\widetilde{B}_{3}$, then as before we know $\Delta_{\Lambda_{1}}$ satisfies the assumptions of Proposition 3.35, and Proposition 8.1 (1) implies that $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ is bowtie free and upward flag. Thus the claim is proved. Similarly, if $\Lambda_{4}$ is the component of $\Lambda\setminus\{s_{4}\}$ containing $\Lambda^{\prime}\setminus\{s_{4}\}$, then the vertex set of $\Delta_{\Lambda_{4},\Lambda_{4}\cap\Lambda^{\prime}}$, endowed with the order induced from $s_{1}<s_{2}<s_{3}$ is a bowtie free and downward flag poset. As $\Lambda$ is a tree, $\Delta_{\Lambda,\Lambda^{\prime}}$ with its vertices endowed with the order induced from $s_{1}<s_{2}<s_{3}<s_{4}$ is a poset. Now Theorem 3.27 implies that $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible.
@@
Case 2: $\Lambda^{\prime}$ is of type $\widetilde{B}_{3}$. Let $a$ be the center vertex and $\{b_{i}\}_{i=1}^{3}$ be leave vertices of $\Lambda^{\prime}$ with $m_{a,b_{1}}=4$. For $1\leq i\leq 3$, let $\Lambda_{i}$ be the component of $\Lambda\setminus\{b_{i}\}$ containing $a$. Now we verify the assumptions of Proposition 3.35 hold for $\Delta_{\Lambda,\Lambda^{\prime}}$. As $\Lambda_{i}$ contains an edge labeled $4$ for $i=2,3$, we know the type of $\Lambda_{i}$ belongs to $\{F_{4},B_{n},\widetilde{C}_{n},\widetilde{B}_{3}\}$. Then Assumption 1 of Proposition 3.35 follows from Proposition 4.2 if $\Lambda_{i}$ is of type $F_{4}$, Proposition 3.22 if $\Lambda_{i}$ is of type $B_{n}$, Corollary 5.2 if $\Lambda_{i}$ is of type $\widetilde{C}_{n}$, and Proposition 8.1 (1) if $\Lambda_{i}$ is of type $\widetilde{B}_{3}$. For Assumption 2 of Proposition 3.35, note that the type of $\Lambda_{1}$ belongs to $\{A_{n},B_{n},D_{4},\widetilde{C}_{n},\widetilde{B}_{3}\}$ (otherwise we will have a contradiction with the fact that the type of $\Lambda_{i}$ belongs to $\{F_{4},B_{n},\widetilde{C}_{n},\widetilde{B}_{3}\}$ for $i=2,3$). Thus Assumption 2 of Proposition 3.35 follows from Theorem 5.6 if $\Lambda_{1}$ is of type $A_{n}$, Proposition 8.1 (2) if $\Lambda_{1}$ is of type $\widetilde{B}_{3}$, Theorem 3.16 if $\Lambda_{1}$ is of type $D_{4}$, and Lemma 11.2 below if $\Lambda_{1}$ is of type $B_{n}$ or $\widetilde{C}_{n}$. ∎

### v3

@@ 11.2 Proof of contractibility
Suppose $\Lambda=\Lambda_{m,n}$ which is the join of $m$ nodes and $n$ nodes. The case that one of $m$ and $n$ is 1 follows from Lemma 11.7. Now we assume $m>1$ and $n>1$. Let $f:\Lambda\to C$ be as in Lemma 11.10. We claim whenever $f$ is a composition of two special foldings $f^{\prime}:\Lambda\to\Lambda^{\prime}$ and $f^{\prime\prime}:\Lambda^{\prime}\to C$ with $\Lambda^{\prime}$ being complete bipartite, then $\Delta_{\Lambda,f^{\prime}}$ is contractible. The lemma follows from this claim by taking $f^{\prime\prime}$ to be identity.
@@
such that $f=f_{1}\circ f_{2}\circ\cdots\circ f_{n}$, each $\Lambda_{i}$ is complete bipartite and $\Lambda_{i+1}$ has exactly one node less than $\Lambda_{i}$. Let $g_{i}=f_{1}\circ\cdots\circ f_{i}$ and $\Delta_{i}=\Delta_{\Lambda,g_{i}}$. The sequence can be arranged such that $\Delta_{\Lambda,f^{\prime}}=\Delta_{i}$ for some $i$. So it suffices to show $\Delta_{i}$ is contractible for any $i$. We set $\Delta_{0}=\Delta_{\Lambda}$.
For each $i$, let $v_{i}\in\Lambda_{i}$ be the unique node such that $f^{-1}_{i}(v_{i})$ has more than one nodes. For each edge in $\Delta_{i}$ whose vertices are of type $f^{-1}_{i+1}(v_{i+1})$, we add a new vertex in this edge which is the midpoint of this edge, and say this new vertex has type $\hat{m}_{i}$. Cut each top dimensional simplex in $\Delta_{i}$ along the new vertex into two simplices, and let the resulting complex by $\Delta^{\prime}_{i}$. Then there is a natural embedding $\iota_{i+1}:\Delta_{i+1}\to\Delta_{i}$ mapping vertices of type $\hat{v}$ to vertices of type $\hat{f}^{-1}_{i+1}(v)$ for $v\in\Lambda_{i+1}\setminus\{v_{i+1}\}$, and mapping vertices of type $\hat{v}_{i+1}$ to vertices of type $\hat{m}_{i}$. The image of $\iota_{i+1}$ is the full subcomplex of $\Delta_{i}$ spanned by vertices whose types are either $\hat{m}_{i}$ or inside $f^{-1}_{i+1}(\Lambda_{i+1}\setminus\{v_{i+1}\})$.
By Lemma 11.10, Theorem 2.19 and [Hua24a, Lemma 10.3], we know $\Delta_{n}$ is contractible. Next we will show $\Delta_{i}$ and $\Delta_{i+1}$ are homotopic equivalent for $10\leq i\leq n-$ 1. By the description of how $\Delta_{i+1}$ sits as a subcomplex of $\Delta^{\prime}_{i}$ via $\iota_{i+1}$ in the previous paragraph, it suffices to show for any $v\in f^{-1}_{i+1}(v_{i+1})$ and any vertex $x\in\Delta_{i}$ of type $\hat{v}$, $\operatorname{lk}(x,\Delta_{i})$ is contractible, as this would imply $\Delta^{\prime}_{i}$ deformation retracts onto $\Delta_{i+1}$. As $\Lambda_{i}$ is complete bipartite, $f^{-1}_{i+1}(v_{i+1})$ is contained in a join factor of $\Lambda^{\prime}_{i}$ of $\Lambda_{i}$. As $f^{-1}_{i+1}(v_{i+1})$ has two elements, we know $\Lambda_{i}\setminus\{v\}$ is connected. Let $\Lambda_{i,v}$ be the unique component of $\Lambda_{i}\setminus\{v\}$. Then $\Lambda_{i,v}$ is also complete bipartite. Let $\Theta=g^{-1}_{i}(\Lambda_{i,v})$. By [Hua24a, Lemma 10.4],
@@
Let $\Lambda$ be a complete bipartite Coxeter diagram. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
@@
Note that when $\Lambda$ has $\leq 4$ nodes, $A_{\Lambda}$ satisfies $K(\pi,1)$-conjecture by Theorem 10.3. Now the theorem follows by induction on the number of nodes in $\Lambda$, using Proposition 11.4, Lemma 11.11 and Theorem 2.5. ∎
@@
Let $\Lambda$ be a tree Coxeter diagram with a collection of open edges $E$ with label $\geq 6$ such that each component of $\Lambda\setminus(\cup_{e\in E}\{e\})$ is either spherical or a star. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
@@
By Theorem 1.8 and [Hua24a, Proposition 9.12], it suffices to show $\Delta_{\Lambda}$ satisfies labeled 4-wheel condition whenever $\Lambda$ is a star. Take a maximal linear subgraph $\Lambda^{\prime}\subset\Lambda$. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is bowtie free - the case when $\Lambda$ has $\leq 3$ nodes follows from Theorem 2.35 and Lemma 11.5, the case when $\Lambda$ has 4 nodes follows from Proposition 11.4 and the case when $\Lambda$ has $\geq 5$ nodes follows from Lemma 11.7. Then Lemma 2.30 implies that $\Delta_{\Lambda}$ satisfies labeled 4-wheel condition. ∎
@@ 12 $K(\pi,1)$ for some higher-dimensional families
Let $\Lambda$ be a connected Coxeter diagram with an induced sub-diagram $\Lambda^{\prime}\subset\Lambda$ such that
@@
    $\Lambda^{\prime}$ is the Coxeter diagram of a 3-dimensional irreducible affine Coxeter group;
@@
    for any node $s\in\Lambda^{\prime}$, each component of $\Lambda\setminus\{s\}$ is either spherical, or has type in $\{\widetilde{A}_{3},\widetilde{B}_{3},\widetilde{C}_{n}\}$.
@@
Let $\mathcal{C}$ be the class of Coxeter diagrams satisfying the assumptions of Proposition 12.1. Let $\Lambda_{0}$ be a 5-cycle with its consecutive edges labeled by $\{3,3,4,3,4\}$. Let $\mathcal{C}_{T}$ be the subclass made of members of $\mathcal{C}$ that are trees. Let $\mathcal{C}^{\prime}$ be the collection of all $\Lambda\in\mathcal{C}$ such that there exists $\Lambda^{\prime}\subset\Lambda$ of type $\widetilde{A}_{3}$ with Assumption 2 satisfied. We first show that element of $\mathcal{C}=\mathcal{C}_{T}\sqcup\mathcal{C}^{\prime}\sqcup\{\Lambda_{0}\}$. Indeed, take $\Lambda\in\mathcal{C}$ and let $\Lambda^{\prime}\subset\Lambda$ be as in Assumption 2. First we consider the case $\Lambda^{\prime}$ has type $\widetilde{C}_{3}$. Let $\{s_{i}\}_{i=1}^{4}$ be consecutive nodes in $\Lambda^{\prime}$. By our assumption and the classification of spherical and Euclidean Coxeter diagrams, for $i=1,4$, the only possibility types for each component of $\Lambda\setminus\{s_{i}\}$ are $\{B_{n},F_{4},\widetilde{B}_{3},\widetilde{C}_{n}\}$. If $\Lambda$ is not a tree, then it must contain an embedded cycle $C$. As each component of $\Lambda\setminus\{s_{i}\}$ is a tree for $i=1,4$, we know $\{s_{1},s_{4}\}\subset C$. Note that $s_{2}\in C$, otherwise $C\cup\overline{s_{3}s_{4}}$ is contained in a component of $\Lambda\setminus\{s_{2}\}$, which is not possible by our assumption (as the label of $\overline{s_{3}s_{4}}$ is $4$). Similarly $s_{3}\in C$. Thus $\Lambda^{\prime}\subset C$. Let $e_{i}$ be the edge of $C$ that is outside $\Lambda^{\prime}$ and contains $s_{i}$ for $i=1,4$. As $\Lambda^{\prime}$ is an induced subgraph of $\Lambda$, $s_{1}\notin e_{4}$ and $s_{4}\notin e_{1}$. Thus $e_{1}\cup\overline{s_{1}s_{2}}\cup\overline{s_{2}s_{3}}$ is contained in a component of $\Lambda\setminus\{s_{4}\}$, and by our assumption, the only possibility of this component is $F_{4}$. Similarly, the component of $\Lambda\setminus\{s_{1}\}$ containing $\overline{s_{2}s_{3}}\cup\overline{s_{3}s_{4}}\cup e_{4}$ is of type $F_{4}$. It follows from Assumption 2 of the proposition that $\Lambda=\Lambda_{0}$. Now we consider the case that $\Lambda^{\prime}$ has type $\widetilde{B}_{3}$. Let $a$ be the center node and $\{b_{i}\}_{i=1}^{3}$ be leave nodes of $\Lambda^{\prime}$ with $m_{a,b_{1}}=4$. For $1\leq i\leq 3$, let $\Lambda_{i}$ be the component of $\Lambda\setminus\{b_{i}\}$ containing $a$. Then for $i=2,3$, $\Lambda_{i}$ contains an edge labeled by 4, hence is a tree by the classification of spherical and Euclidean Coxeter diagrams. If $\Lambda_{1}$ is a tree, then $\Lambda$ is a tree. Otherwise $\Lambda_{1}$ is of type $\widetilde{A}_{3}$. Then for $i=2,3$, $\Lambda_{i}$ contains a path of length 3, with its edges labeled by $4,3,3$. Thus $\Lambda_{i}$ is of type $B_{n}$ or $\widetilde{C}_{n}$ for $i=2,3$. However, only type $B_{n}$ and $n=4$ is possible, otherwise $\Lambda_{1}$ is not of type $\widetilde{A}_{3}$. This implies that $\Lambda$ is obtained from a diagram of type $\widetilde{A}_{3}$ by adding an extra edge of label $4$. Thus $\Lambda\in\mathcal{C}^{\prime}$.
@@
Next we show $\mathcal{C}^{\prime}$ satisfy the assumptions of [Hua24a, Corollary 7.3] with $\mathcal{C}_{2}=\mathcal{C}^{\prime}$, and $\mathcal{C}_{1}$ be the class containing only the diagram of type $\widetilde{A}_{3}$, hence each element in $\mathcal{C}^{\prime}$ satisfies the $K(\pi,1)$-conjecture. Note that Assumptions $2,3,4$ of [Hua24a, Corollary 7.3] are clear. It remains to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible. For node $s\in\Lambda^{\prime}$, Let $\Lambda_{s}$ be the component of $\Lambda\setminus\{s\}$ containing the rest of the nodes of $\Lambda^{\prime}$. If $\Lambda_{s}$ is not a tree for some $s$, then $\Lambda_{s}$ is of type $\widetilde{A}_{3}$. Then $\Lambda$ is a complete bipartite graph $K_{2,3}$. By [Hua24a, Theorem 10.7 and Corollary 10.10], $\Delta_{\Lambda}$ is contractible. Hence $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible by Lemma 2.4. Now we assume $\Lambda_{s}$ is a tree for each $s$. By Theorem 2.19 and Lemma 2.2, it suffices to show $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. However, this follows from Theorem 2.35 if $\Lambda_{s}$ is spherical, and Corollary 5.2 if $\Lambda_{s}$ is of type $\widetilde{C}_{n}$. If $\Lambda_{s}$ is of type $\widetilde{B}_{3}$, then by Theorem 2.37 and Theorem 5.6, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 8.3. Thus by Proposition 9.1 $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free.
@@
As the $K(\pi,1)$-conjecture for Artin group with diagram $\Lambda_{0}$ follows from [Hua24a, Theorem 10.9], it remains to show diagrams in $\mathcal{C}_{T}$ satisfies $K(\pi,1)$-conjecture. We will show $\mathcal{C}_{T}$ satisfy the assumptions of [Hua24a, Corollary 7.3] with $\mathcal{C}_{2}=\mathcal{C}_{T}$ and $\mathcal{C}_{1}=\{\widetilde{C}_{3},\widetilde{B}_{3}\}$. Again Assumptions $2,3,4$ of [Hua24a, Corollary 7.3] are clear, and it suffices to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible with $\Lambda^{\prime}\in\mathcal{C}_{1}$ and $\Lambda\in\mathcal{C}_{2}$.
@@
Case 1: $\Lambda^{\prime}$ is of type $\widetilde{C}_{3}$. Let consecutive nodes of $\Lambda^{\prime}$ be $\{s_{i}\}_{i=1}^{4}$. Let $\Lambda_{1}$ be the component of $\Lambda\setminus\{s_{1}\}$ containing $\Lambda^{\prime}\setminus\{s_{1}\}$. By Assumption 2, $\Lambda_{1}$ is either spherical or irreducible 3-dimensional Euclidean. We claim the vertex set of $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$, endowed with the order induced from $s_{2}<s_{3}<s_{4}$ is a bowtie free and upward flag poset. As $\Lambda_{1}$ is a tree, $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ is indeed a poset. If $\Lambda_{1}$ is spherical, then either $\Lambda_{1}$ is of type $B_{n}$ or type $F_{4}$. Then $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ being bowtie free and upward flag follows from Theorem 2.35, Theorem 2.37 and Proposition 3.2. If $\Lambda_{1}$ is of type $\widetilde{C}_{n}$, then the claim follows from Corollary 5.2. If $\Lambda_{1}$ is of type $\widetilde{B}_{3}$, then as before we know $\Delta_{\Lambda_{1}}$ satisfies the assumptions of Proposition 8.3, and Proposition 9.1 (1) implies that $\Delta_{\Lambda_{1},\Lambda_{1}\cap\Lambda^{\prime}}$ is bowtie free and upward flag. Thus the claim is proved. Similarly, if $\Lambda_{4}$ is the component of $\Lambda\setminus\{s_{4}\}$ containing $\Lambda^{\prime}\setminus\{s_{4}\}$, then the vertex set of $\Delta_{\Lambda_{4},\Lambda_{4}\cap\Lambda^{\prime}}$, endowed with the order induced from $s_{1}<s_{2}<s_{3}$ is a bowtie free and downward flag poset. As $\Lambda$ is a tree, $\Delta_{\Lambda,\Lambda^{\prime}}$ with its vertices endowed with the order induced from $s_{1}<s_{2}<s_{3}<s_{4}$ is a poset. Now Theorem 2.18 implies that $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible.
@@
Case 2: $\Lambda^{\prime}$ is of type $\widetilde{B}_{3}$. Let $a$ be the center node and $\{b_{i}\}_{i=1}^{3}$ be leave nodes of $\Lambda^{\prime}$ with $m_{a,b_{1}}=4$. For $1\leq i\leq 3$, let $\Lambda_{i}$ be the component of $\Lambda\setminus\{b_{i}\}$ containing $a$. Now we verify the assumptions of Proposition 8.3 hold for $\Delta_{\Lambda,\Lambda^{\prime}}$. As $\Lambda_{i}$ contains an edge labeled $4$ for $i=2,3$, we know the type of $\Lambda_{i}$ belongs to $\{F_{4},B_{n},\widetilde{C}_{n},\widetilde{B}_{3}\}$. Then Assumption 1 of Proposition 8.3 follows from Proposition 3.2 if $\Lambda_{i}$ is of type $F_{4}$, Proposition 2.37 if $\Lambda_{i}$ is of type $B_{n}$, Corollary 5.2 if $\Lambda_{i}$ is of type $\widetilde{C}_{n}$, and Proposition 9.1 (1) if $\Lambda_{i}$ is of type $\widetilde{B}_{3}$. For Assumption 2 of Proposition 8.3, note that the type of $\Lambda_{1}$ belongs to $\{A_{n},B_{n},D_{4},\widetilde{C}_{n},\widetilde{B}_{3}\}$ (otherwise we will have a contradiction with the fact that the type of $\Lambda_{i}$ belongs to $\{F_{4},B_{n},\widetilde{C}_{n},\widetilde{B}_{3}\}$ for $i=2,3$). Thus Assumption 2 of Proposition 8.3 follows from Theorem 5.6 if $\Lambda_{1}$ is of type $A_{n}$, Proposition 9.1 (2) if $\Lambda_{1}$ is of type $\widetilde{B}_{3}$, Theorem 2.38 if $\Lambda_{1}$ is of type $D_{4}$, and Lemma 12.2 below if $\Lambda_{1}$ is of type $B_{n}$ or $\widetilde{C}_{n}$. ∎